OpenStudy (anonymous):

The sum of two numbers is 86. One of the numbers is five less than three times the other. What are the two numbers?

OpenStudy (anonymous):

x + (3x - 5) = 86 4x - 5 = 86 4x = 91 x = 22.75 the first number is 22.75 86 - 22.75 = 63.25 the second number is 63.25